OpenStudy (anonymous):

Translate into an equation: y varies jointly with b and c. Choose one answer. a. y = bc/k b. y = kb/c c. y = k/bc d. y = kbc

OpenStudy (anonymous):

y=kbc d

OpenStudy (anonymous):

a

myininaya (myininaya):

d

OpenStudy (saifoo.khan):

d. y = kbc

OpenStudy (anonymous):

a and d are the same if k is a constant

OpenStudy (anonymous):

the constant is never the denominator y=kx directly y=k/x or xy=k is inverse

myininaya (myininaya):

what jzile says is true though a=d

OpenStudy (anonymous):

peaceout:this is this same as your last question....referr to that for the answer

myininaya (myininaya):

i will give you an example of why a and d are the same say i ask you to find k of y=kbc if b=2 c=3 and y=54, then you have: 54=k(2)(3)=> k=9 so you have y=9bc now say i give you the same b and c, b=2, b=3 and same y,y=54 and find k if y=bc/k 54=(2)(3)/k 54k=6 so k=6/54=1/9 so we have y=bc/(1/9)=9bc the equations are the same

OpenStudy (anonymous):

but if the question asks for the constant of variation then 9 and 1/9 are different answers

myininaya (myininaya):

no but 9 and 1/(1/9) are the same

OpenStudy (anonymous):

right but if k is the denominator k is not 1/(1/9) it is just 1/9 and some algebra books just ask for the value of k. I agree with you that a constant is just a constant but with Algebra 1 or even algebra 2, it may be considered an incorrect answer.